# Parallel Diffusion Solver via Residual Dirichlet Policy Optimization

### v1

@@ I Introduction
Diffusion models (DMs) [sohl2015deep, ho2020denoising, rombach2022high] have emerged as a leading paradigm in generative modeling, delivering state-of-the-art performance across image synthesis [rombach2022high, saharia2022photorealistic, Lei_2025_CVPR] and video generation [blattmann2023align, ho2022video, zhou2025streamingdragorientedinteractivevideo, zhao2025realtimemotioncontrollableautoregressivevideo]. These models generate data by iteratively refining noisy inputs through a sequential denoising process. While this mechanism produces high-fidelity outputs, the requirement for multi-step sequential evaluation introduces substantial latency, rendering real-time sampling inefficient.
@@
Figure 2: Computation graphs of various ODE solvers. (a) DDIM solver [songdenoising] (Euler’s method) adopts the rectangle rule that uses the gradient at the start point: ${\mathbf{d}}_{t_{n+1}}=\bm{\epsilon}_{\theta}({\mathbf{x}}_{t_{n+1}},t_{n+1})$. disclose EDM solver [karras2022elucidating] (Heun’s method) uses the trapezoidal rule that averages the gradients of both the start and the end timesteps, i.e., ${\mathbf{d}}_{t_{n+1}}=\bm{\epsilon}_{\theta}({\mathbf{x}}_{t_{n+1}},t_{n+1})$ and ${\mathbf{d}}_{t_{n}}^{\prime}=\bm{\epsilon}_{\theta}({\mathbf{x}}_{t_{n}}^{\prime},t_{n})$, where ${\mathbf{x}}_{t_{n}}^{\prime}$ is the additional evaluation given by Euler’s method. (c) AMED solver [zhou2024fast] optimizes a small network $g_{\phi}(\cdot)$ to output an intermediate timestep $s_{n}\in(t_{n},t_{n+1})$ to compute the gradient: ${\mathbf{d}}_{s_{n}}=\bm{\epsilon}_{\theta}({\mathbf{x}}_{s_{n}},s_{n})$. Since AMED introduces a network in sequential computation, its latency is slightly higher than that of other solvers, as shown in fig. 1. (d) Our EPD-Solver leverage $K$ parallel gradients to achieve more accurate integral approximation. We optimize $K$ intermediate timesteps $\tau_{n}^{1},\dots,\tau_{n}^{K}$, compute their gradients ${\mathbf{d}}_{\tau_{n}^{1}},\dots,{\mathbf{d}}_{\tau_{n}^{K}}$, and combine them via a simplex-weighted sum.
@@
To mitigate this inefficiency, recent research has focused on accelerating the sampling process through various approaches, such as solver-based, distillation-based and parallelism-based methods. Solver-based methods aim to develop fast numerical solvers to reduce the number of sampling steps [songdenoising, karras2022elucidating, lu2022dpm, lu2022dpm_plus, liupseudo, zhangfast, zhao2024unipc, zhou2024fast, kim2024distilling, watson2021learning]. However, aggressive step reduction often leads to significant truncation errors, causing quality degradation at low function evaluations (NFEs). Distillation-based methods establish a direct mapping between noise and data [zhou2025simple, luhman2021knowledge, liuflow, berthelot2023tract, salimansprogressive, meng2023distillation, song2023consistency, luo2023latent, kimconsistency], achieving extreme acceleration (e.g., one-step generation). Yet, they incur high training costs and lack the flexibility to trade speed for quality. Parallelism-based methods [shih2023parallel, li2024faster, li2024distrifusion, chenasyncdiff] attempt to trade computation for speed, but this direction remains under-explored for quality enhancement.
@@
In this paper, we seek to combine the strengths of these approaches by investigating solver-based methods under low-latency constraints. We propose the Ensemble Parallel Direction (EPD) solver, a novel approach that leverages extra parallel computation to minimize truncation errors in each ODE step without increasing wall-clock time. Unlike standard solvers that rely on a single gradient evaluation (e.g., DDIM [songdenoising]) or sequential multi-point estimates (e.g., EDM [karras2022elucidating] and AMED [zhangfast]), our method concurrently evaluates gradients at multiple learned intermediate timesteps within a single integration interval (see Figure 2). By aggregating these parallel gradient estimates via a weighted combination, we achieve a significantly more accurate approximation of the integral direction, which is theoretically grounded in the mean value theorem for vector-valued functions [mcleod1965mean]. Crucially, since these gradient computations are independent of each other, they can be efficiently parallelized on modern hardware. This allows EPD-Solver to enhance sampling fidelity with negligible latency overhead. As in Figure 1, EPD-Solver consistently achieves better FID scores than existing ODE solvers at comparable inference latencies on CIFAR [krizhevsky2009learning].
@@
We adopt a two-stage optimization framework to decide the intermediate evaluation timesteps and their combining weights. (1) In the first stage, we distill a few-step EPD sampler by optimizing its learnable parameters to approximate the trajectories generated by a high-NFE teacher solver. However, in the extremely low-step regime, distillation alone is insufficient. It not only struggles to learn an accurate mapping from noise to teacher trajectories [tong2024learning] but also falls short in aligning with human perceptual preferences. For large-scale text-to-image (T2I) diffusion models, human preference is better characterized by semantic and perceptual alignment rather than strict trajectory consistency. Motivated by this, (2) in the second stage, we perform Residual Dirichlet Policy Optimization (RDPO), where the solver is reparameterized as a stochastic policy initialized using the parameters distilled in Stage 1. This parameterization induces a structured, simplex-constrained policy space that supports stable and efficient optimization of human-aligned rewards via a PPO [schulman2017proximalpolicyoptimizationalgorithms] variant. Moreover, our framework is lightweight and plug-and-play: it optimizes only a few solver parameters while freezing the backbone, thereby reducing tuning cost, improving RL stability, and preserving generation robustness; the resulting solver can also be seamlessly integrated into existing ODE samplers as EPD-Plugin.
@@
We evaluate EPD-Solver across a diverse set of image generation models spanning resolutions from 32 to 1024, including standard unconditional and class-conditional benchmarks (CIFAR-10 [krizhevsky2009learning], FFHQ [karras2019style], ImageNet [russakovsky2015imagenet], LSUN Bedroom [yu2015lsun]) and large-scale T2I models (Stable Diffusion v1.5 [rombach2022high], SD3-Medium [sd3]). Empirical results confirm that incorporating parallel gradients significantly reduces truncation errors and consistently outperforms prior learning-based solvers. At 5 NFE, EPD-Solver achieves FIDs of 4.47 (CIFAR-10), 7.97 (FFHQ), 8.17 (ImageNet), and 8.26 (LSUN Bedroom), notably surpassing AMED-Solver [zhou2024fast], which yields 13.20 on LSUN Bedroom. For T2I models, our residual Dirichlet policy based solver achieves strong alignment with high efficiency: at just 20 NFEs, EPD-Solver attains an HPSv2.1 score of 0.2482 and an ImageReward of 3.1121, surpassing 50-step baselines like iPNDM while cutting inference cost by 60%. Our contributions are summarized as follows:
@@
Our preliminary results was published in ICCV 2025 [zhu2025distilling]. The source code and checkpoints are available in https://github.com/BeierZhu/EPD.
@@ II-A Sampling Acceleration Methods
High latency in the sampling process is a major drawback of DMs compared to other generative models [goodfellow2014generative, kingma2013auto]. Prior acceleration efforts mainly fall into the following categories:
@@
Distillation-based methods. These methods accelerate diffusion models by re-training or fine-tuning the entire DM. One category is trajectory distillation, which trains a student model to imitate the teacher’s trajectory with fewer steps [zhou2025simple]. This process can be achieved through offline distillation [luhman2021knowledge, liuflow], which requires constructing a dataset sampled from teacher models, or online distillation, which progressively reduces sampling steps in a multi-stage manner [berthelot2023tract, salimansprogressive, meng2023distillation]. Another line of research is consistency distillation, where the denoising outputs along the sampling trajectory are enforced to remain consistent [song2023consistency, luo2023latent, kimconsistency]. Apart from distilling noise-image pairs, distribution matching methods match real and reconstructed samples at the distribution level [pooledreamfusion, wang2023prolificdreamer, sauer2024adversarial, yin2024one]. Despite significantly enhancing quality, these approaches incur high training costs and require carefully designed training procedures.
@@
Solver-based methods. Beyond fine-tuning DMs, fast ODE solvers have been extensively studied. Training-free methods include Euler’s method [songdenoising], Heun’s method [karras2022elucidating], Taylor expansion-based solvers (DPM-Solver [lu2022dpm], DPM-Solver++ [lu2022dpm_plus]), multi-step methods (PNDM [liupseudo], iPNDM [zhangfast]), and predictor-corrector frameworks (UniPC [zhao2024unipc]). Some solvers require additional training, e.g., AMED-Solver [zhou2024fast] , D-ODE [kim2024distilling], and DDSS [watson2021learning], AdaSDE [wang2025adaptivestochasticcoefficientsaccelerating]. Recent work optimizes timestep schedules, with notable studies including LD3 [tong2024learning], AYS [sabour2024align], GITS [chen2024trajectory], and DMN [xue2024accelerating]. Though EPD-Solver falls into this category, we optimize solver parameters via distillation to achieve high-quality, low-latency generation through parallelism. With minimal learnable parameters, training remains highly efficient.
@@
Parallelism-based methods. While promising, parallelism remains an underexplored approach for accelerating diffusion models. ParaDiGMS [shih2023parallel] leverages Picard iteration for parallel sampling but struggles to maintain consistency with original outputs. Faster Diffusion [li2024faster] performs decoder computation in parallel by omitting encoder computation at some adjacent timesteps, but this compromises image quality. Distrifusion [li2024distrifusion] divides high-resolution images into patches and performs parallel inference on each patch. AsyncDiff [chenasyncdiff] implements model parallelism through asynchronous denoising. Unlike prior methods that focus on reducing latency, our EPD-Solver leverages parallel gradients to enhance image quality without incurring notable latency.
@@
Beyond the above three categories, cache-based acceleration methods [ma2023deepcache, zhao2024dynamic, zou2024accelerating, Liu_2025_CVPR, liu2024timestep] aim to reduce per-step computational overhead by exploiting temporal redundancy during diffusion sampling. Specifically, these approaches reuse feature maps or intermediate states from preceding steps to bypass redundant computations. While effective in reducing wall-clock latency, they do not directly address discretization errors inherent in the numerical integration of the sampling trajectory and are largely orthogonal to our approach.
@@ II-B Reinforcement Learning from Human Feedback
RL-based approaches for aligning pretrained text-to-image (T2I) diffusion models with human preferences can be broadly categorized into supervised-style objectives and policy-gradient–based RL. Supervised-style methods formulate alignment as weighted maximum likelihood or preference-matching, directly shaping the data distribution using scalar rewards or pairwise preference signals without explicit policy gradients [clark2024directlyfinetuningdiffusionmodels, xu2023imagereward, peng2019advantageweightedregressionsimplescalable, lee2023aligningtexttoimagemodelsusing, Wallace_2024_CVPR, dong2023raftrewardrankedfinetuning, zheng2025diffusionnftonlinediffusionreinforcement, zhao2025unsupervised]. In contrast, policy-gradient–based methods treat the T2I model as a stochastic policy and explicitly optimize expected rewards via RL updates, typically in a PPO-style framework [black2024trainingdiffusionmodelsreinforcement, NEURIPS2023_fc65fab8, gupta2025simpleeffectivereinforcementlearning, Miao_2024_CVPR, zhao2025scoreactionfinetuningdiffusion, liu2025flowgrpotrainingflowmatching, li2025mixgrpounlockingflowbasedgrpo, xue2025dancegrpounleashinggrpovisual]. Different from existing RL-based alignment methods that optimize the DM itself, we perform RL at the solver level by learning a residual Dirichlet policy around a distilled base solver, enabling both parameter-efficient and robust preference alignment.
@@ III-A Background
Diffusion models gradually inject noise into data via a forward noising process and generate samples by learning a reversed denoising process, initialized with Gaussian noise. Let ${\mathbf{x}}\sim p_{\text{data}}({\mathbf{x}})$ denote the $d$-dimensional data and $p({\mathbf{x}};\sigma)$ the data distribution with Gaussian noise of variance $\sigma^{2}$ injected. The forward process is controlled by a noise schedule defined by the time scaling $s(t)$ and the noise level $\sigma(t)$ at time $t$. In particular, ${\mathbf{x}}=s(t)\hat{{\mathbf{x}}}_{t}$, where $\hat{{\mathbf{x}}}_{t}\sim p({\mathbf{x}};\sigma(t))$. Such forward process can be formulated by a SDE [karras2022elucidating]:
@@
where ${\mathbf{w}}\in{\mathbb{R}}^{d}$ denotes Wiener process. In this paper, we adopt the framework of EDM [karras2022elucidating] by setting $\sigma(t)=t$ and $s(t)=1$. Generation is then performed with the reverse of eq. 1. Notably, there exists the probability flow ODE:
@@
The noise-prediction model $\bm{\epsilon}_{\theta}({\mathbf{x}},t)$ is trained by minimizing the $\ell_{2}^{2}$ loss with a weighting function $\lambda(t)$ [karras2022elucidating, song2021scorebased]:
@@ III-B Motivation and Analysis
Various ODE solvers have been proposed to approximate the integral. At a high level, these solvers leverage one or several points to compute gradients, which are then used to estimate the integral. Let $I$ denote the integral $I=\int_{t_{n+1}}^{t_{n}}\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t){\mathrm{d}}t$ and $h_{n}$ denote the step length $h_{n}=t_{n}-t_{n+1}$. For instance, DDIM [songdenoising] (Euler’s method) adopts the rectangle rule that uses the gradient at the start point:
@@
EDM [karras2022elucidating] considers the trapezoidal rule that averages the gradients of both the start and end points.
@@
where ${\mathbf{x}}_{t_{n}}^{\prime}$ is the additional evaluation point given by Euler’s method, i.e., ${\mathbf{x}}_{t_{n}}^{\prime}={\mathbf{x}}_{t_{n+1}}+h_{n}\bm{\epsilon}_{\theta}({\mathbf{x}}_{t_{n+1}},t_{n+1})$. AMED-Solver [zhou2024fast] optimizes a small network to output an intermediate timestep $s_{n}\in(t_{n},t_{n+1})$ to compute the gradient:
@@
([mcleod1965mean]) When $f$ has values in an $n$-dimensional vector space and is continuous on the closed interval $[a,b]$ and differentiable on the open interval $(a,b)$, we have
@@
Figure 3: Cumulative explained variance ratio of sampling trajectories using DMs from EDM2 [Karras2024edm2]. We analyze the trajectory’s orthogonal complement, i.e., the residuals after removing the linear component connecting $\mathbf{x}_{t_{T}}$ and $\mathbf{x}_{t_{0}}$. The rapid saturation at the two principle components (capturing $97\%$ of the residual variance) indicates that the trajectory occurs almost within a single 2D plane.

### v2

@@ I Introduction
Diffusion models (DMs) [54, 13, 46] have emerged as a leading paradigm in generative modeling, delivering state-of-the-art performance across image synthesis [46, 49, 23] and video generation [3, 14, 78, 73]. These models generate data by iteratively refining noisy inputs through a sequential denoising process. While this mechanism produces high-fidelity outputs, the requirement for multi-step sequential evaluation introduces substantial latency, rendering real-time sampling inefficient.
@@
Figure 2: Computation graphs of various ODE solvers. (a) DDIM solver [55] (Euler’s method) adopts the rectangle rule that uses the gradient at the start point: ${\mathbf{d}}_{t_{n+1}}=\bm{\epsilon}_{\theta}({\mathbf{x}}_{t_{n+1}},t_{n+1})$. disclose EDM solver [15] (Heun’s method) uses the trapezoidal rule that averages the gradients of both the start and the end timesteps, i.e., ${\mathbf{d}}_{t_{n+1}}=\bm{\epsilon}_{\theta}({\mathbf{x}}_{t_{n+1}},t_{n+1})$ and ${\mathbf{d}}_{t_{n}}^{\prime}=\bm{\epsilon}_{\theta}({\mathbf{x}}_{t_{n}}^{\prime},t_{n})$, where ${\mathbf{x}}_{t_{n}}^{\prime}$ is the additional evaluation given by Euler’s method. (c) AMED solver [80] optimizes a small network $g_{\phi}(\cdot)$ to output an intermediate timestep $s_{n}\in(t_{n},t_{n+1})$ to compute the gradient: ${\mathbf{d}}_{s_{n}}=\bm{\epsilon}_{\theta}({\mathbf{x}}_{s_{n}},s_{n})$. Since AMED introduces a network in sequential computation, its latency is slightly higher than that of other solvers, as shown in fig. 1. (d) Our EPD-Solver leverage $K$ parallel gradients to achieve more accurate integral approximation. We optimize $K$ intermediate timesteps $\tau_{n}^{1},\dots,\tau_{n}^{K}$, compute their gradients ${\mathbf{d}}_{\tau_{n}^{1}},\dots,{\mathbf{d}}_{\tau_{n}^{K}}$, and combine them via a simplex-weighted sum.
@@
To mitigate this inefficiency, recent research has focused on accelerating the sampling process through various approaches, such as solver-based, distillation-based and parallelism-based methods. Solver-based methods aim to develop fast numerical solvers to reduce the number of sampling steps [55, 15, 34, 35, 32, 71, 76, 80, 19, 63]. However, aggressive step reduction often leads to significant truncation errors, causing quality degradation at low function evaluations (NFEs). Distillation-based methods establish a direct mapping between noise and data [79, 36, 33, 1, 50, 40, 56, 37, 18], achieving extreme acceleration (e.g., one-step generation). Yet, they incur high training costs and lack the flexibility to trade speed for quality. Parallelism-based methods [53, 27, 26, 5] attempt to trade computation for speed, but this direction remains under-explored for quality enhancement.
@@
In this paper, we seek to combine the strengths of these approaches by investigating solver-based methods under low-latency constraints. We propose the Ensemble Parallel Direction (EPD) solver, a novel approach that leverages extra parallel computation to minimize truncation errors in each ODE step without increasing wall-clock time. Unlike standard solvers that rely on a single gradient evaluation (e.g., DDIM [55]) or sequential multi-point estimates (e.g., EDM [15] and AMED [71]), our method concurrently evaluates gradients at multiple learned intermediate timesteps within a single integration interval (see Figure 2). By aggregating these parallel gradient estimates via a weighted combination, we achieve a significantly more accurate approximation of the integral direction, which is theoretically grounded in the mean value theorem for vector-valued functions [39]. Crucially, since these gradient computations are independent of each other, they can be efficiently parallelized on modern hardware. This allows EPD-Solver to enhance sampling fidelity with negligible latency overhead. As in Figure 1, EPD-Solver consistently achieves better FID scores than existing ODE solvers at comparable inference latencies on CIFAR [22].
@@
We adopt a two-stage optimization framework to decide the intermediate evaluation timesteps and their combining weights. (1) In the first stage, we distill a few-step EPD sampler by optimizing its learnable parameters to approximate the trajectories generated by a high-NFE teacher solver. However, in the extremely low-step regime, distillation alone is insufficient. It not only struggles to learn an accurate mapping from noise to teacher trajectories [59] but also falls short in aligning with human perceptual preferences. For large-scale text-to-image (T2I) diffusion models, human preference is better characterized by semantic and perceptual alignment rather than strict trajectory consistency. Motivated by this, (2) in the second stage, we perform Residual Dirichlet Policy Optimization (RDPO), where the solver is reparameterized as a stochastic policy initialized using the parameters distilled in Stage 1. This parameterization induces a structured, simplex-constrained policy space that supports stable and efficient optimization of human-aligned rewards via a PPO [52] variant. Moreover, our framework is lightweight and plug-and-play: it optimizes only a few solver parameters while freezing the backbone, thereby reducing tuning cost, improving RL stability, and preserving generation robustness; the resulting solver can also be seamlessly integrated into existing ODE samplers as EPD-Plugin.
@@
We evaluate EPD-Solver across a diverse set of image generation models spanning resolutions from 32 to 1024, including standard unconditional and class-conditional benchmarks (CIFAR-10 [22], FFHQ [17], ImageNet [47], LSUN Bedroom [70]) and large-scale T2I models (Stable Diffusion v1.5 [46], SD3-Medium [9]). Empirical results confirm that incorporating parallel gradients significantly reduces truncation errors and consistently outperforms prior learning-based solvers. At 5 NFE, EPD-Solver achieves FIDs of 4.47 (CIFAR-10), 7.97 (FFHQ), 8.17 (ImageNet), and 8.26 (LSUN Bedroom), notably surpassing AMED-Solver [80], which yields 13.20 on LSUN Bedroom. For T2I models, our residual Dirichlet policy based solver achieves strong alignment with high efficiency: at just 20 NFEs, EPD-Solver attains an HPSv2.1 score of 0.2482 and an ImageReward of 0.0121, surpassing 50-step baselines like iPNDM while cutting inference cost by 60%. Our contributions are summarized as follows:
@@
Our preliminary results was published in ICCV 2025 [81]. The source code and checkpoints are available in https://github.com/BeierZhu/EPD.
@@ II-A Sampling Acceleration Methods
High latency in the sampling process is a major drawback of DMs compared to other generative models [11, 20]. Prior acceleration efforts mainly fall into the following categories:
@@
Distillation-based methods. These methods accelerate diffusion models by re-training or fine-tuning the entire DM. One category is trajectory distillation, which trains a student model to imitate the teacher’s trajectory with fewer steps [79]. This process can be achieved through offline distillation [36, 33], which requires constructing a dataset sampled from teacher models, or online distillation, which progressively reduces sampling steps in a multi-stage manner [1, 50, 40]. Another line of research is consistency distillation, where the denoising outputs along the sampling trajectory are enforced to remain consistent [56, 37, 18]. Apart from distilling noise-image pairs, distribution matching methods match real and reconstructed samples at the distribution level [44, 62, 51, 69]. Despite significantly enhancing quality, these approaches incur high training costs and require carefully designed training procedures.
@@
Solver-based methods. Beyond fine-tuning DMs, fast ODE solvers have been extensively studied. Training-free methods include Euler’s method [55], Heun’s method [15], Taylor expansion-based solvers (DPM-Solver [34], DPM-Solver++ [35]), multi-step methods (PNDM [32], iPNDM [71]), and predictor-corrector frameworks (UniPC [76]). Some solvers require additional training, e.g., AMED-Solver [80] , D-ODE [19], and DDSS [63], AdaSDE [61]. Recent work optimizes timestep schedules, with notable studies including LD3 [59], AYS [48], GITS [4], and DMN [66]. Though EPD-Solver falls into this category, we optimize solver parameters via distillation to achieve high-quality, low-latency generation through parallelism. With minimal learnable parameters, training remains highly efficient.
@@
Parallelism-based methods. While promising, parallelism remains an underexplored approach for accelerating diffusion models. ParaDiGMS [53] leverages Picard iteration for parallel sampling but struggles to maintain consistency with original outputs. Faster Diffusion [27] performs decoder computation in parallel by omitting encoder computation at some adjacent timesteps, but this compromises image quality. Distrifusion [26] divides high-resolution images into patches and performs parallel inference on each patch. AsyncDiff [5] implements model parallelism through asynchronous denoising. Unlike prior methods that focus on reducing latency, our EPD-Solver leverages parallel gradients to enhance image quality without incurring notable latency.
@@
Beyond the above three categories, cache-based acceleration methods [38, 75, 82, 31, 29] aim to reduce per-step computational overhead by exploiting temporal redundancy during diffusion sampling. Specifically, these approaches reuse feature maps or intermediate states from preceding steps to bypass redundant computations. While effective in reducing wall-clock latency, they do not directly address discretization errors inherent in the numerical integration of the sampling trajectory and are largely orthogonal to our approach.
@@ II-B Reinforcement Learning from Human Feedback
RL-based approaches for aligning pretrained text-to-image (T2I) diffusion models with human preferences can be broadly categorized into supervised-style objectives and policy-gradient–based RL. Supervised-style methods formulate alignment as weighted maximum likelihood or preference-matching, directly shaping the data distribution using scalar rewards or pairwise preference signals without explicit policy gradients [6, 65, 43, 68, 60, 8, 77, 74]. In contrast, policy-gradient–based methods treat the T2I model as a stochastic policy and explicitly optimize expected rewards via RL updates, typically in a PPO-style framework [2, 10, 12, 41, 72, 30, 24, 67]. Different from existing RL-based alignment methods that optimize the DM itself, we perform RL at the solver level by learning a residual Dirichlet policy around a distilled base solver, enabling both parameter-efficient and robust preference alignment.
@@ III-A Background
Diffusion models gradually inject noise into data via a forward noising process and generate samples by learning a reversed denoising process, initialized with Gaussian noise. Let ${\mathbf{x}}\sim p_{\text{data}}({\mathbf{x}})$ denote the $d$-dimensional data and $p({\mathbf{x}};\sigma)$ the data distribution with Gaussian noise of variance $\sigma^{2}$ injected. The forward process is controlled by a noise schedule defined by the time scaling $s(t)$ and the noise level $\sigma(t)$ at time $t$. In particular, ${\mathbf{x}}=s(t)\hat{{\mathbf{x}}}_{t}$, where $\hat{{\mathbf{x}}}_{t}\sim p({\mathbf{x}};\sigma(t))$. Such forward process can be formulated by a SDE [15]:
@@
where ${\mathbf{w}}\in{\mathbb{R}}^{d}$ denotes Wiener process. In this paper, we adopt the framework of EDM [15] by setting $\sigma(t)=t$ and $s(t)=1$. Generation is then performed with the reverse of eq. 1. Notably, there exists the probability flow ODE:
@@
The noise-prediction model $\bm{\epsilon}_{\theta}({\mathbf{x}},t)$ is trained by minimizing the $\ell_{2}^{2}$ loss with a weighting function $\lambda(t)$ [15, 57]:
@@ III-B Motivation and Analysis
Various ODE solvers have been proposed to approximate the integral. At a high level, these solvers leverage one or several points to compute gradients, which are then used to estimate the integral. Let $I$ denote the integral $I=\int_{t_{n+1}}^{t_{n}}\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t){\mathrm{d}}t$ and $h_{n}$ denote the step length $h_{n}=t_{n}-t_{n+1}$. For instance, DDIM [55] (Euler’s method) adopts the rectangle rule that uses the gradient at the start point:
@@
EDM [15] considers the trapezoidal rule that averages the gradients of both the start and end points.
@@
where ${\mathbf{x}}_{t_{n}}^{\prime}$ is the additional evaluation point given by Euler’s method, i.e., ${\mathbf{x}}_{t_{n}}^{\prime}={\mathbf{x}}_{t_{n+1}}+h_{n}\bm{\epsilon}_{\theta}({\mathbf{x}}_{t_{n+1}},t_{n+1})$. AMED-Solver [80] optimizes a small network to output an intermediate timestep $s_{n}\in(t_{n},t_{n+1})$ to compute the gradient:
@@
([39]) When $f$ has values in an $n$-dimensional vector space and is continuous on the closed interval $[a,b]$ and differentiable on the open interval $(a,b)$, we have
@@
Figure 3: Cumulative explained variance ratio of sampling trajectories using DMs from EDM2 [16]. We analyze the trajectory’s orthogonal complement, i.e., the residuals after removing the linear component connecting $\mathbf{x}_{t_{T}}$ and $\mathbf{x}_{t_{0}}$. The rapid saturation at the two principle components (capturing $97\%$ of the residual variance) indicates that the trajectory occurs almost within a single 2D plane.